OpenStudy (anonymous):

(1 pt) Let f(x)=(1/x-3) and g(x)=(1/x-7). Then the domain of f o g is equal to all reals except for two values, a and b with a

OpenStudy (anonymous):

can you simplify f(g(x)) = ???

OpenStudy (anonymous):

im not sure how to start the problem

OpenStudy (anonymous):

f(x)=(1/x-3) and g(x)=(1/x-7) \(\large f \circ g=f(g(x)=f(\frac{1}{x-7})=\frac{1}{(\frac{1}{x-7})-7} \) simplify this....

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

7(x+7)?